Nine years ago, Kimo was twice as old as Jennifer was then. In four years, Jennifer will be the same age that Kimo is now. How old are they now?
Answer by lwsshak3(11628)

Nine years ago, Kimo was twice as old as Jennifer was then. In four years, Jennifer will be the same age that Kimo is now. How old are they now?
***
9 yrs ago:
let x=Jennifer's age
2x=Kimo's age
..
now:
x+9=Jennifer's age
2x+9=Kimo's age
..
In 4 yrs:
x+9+4=x+13=Jennifer's age
2x+9+4=2x+13=Kimo's age
x+13=2x+9 ( In four years, Jennifer will be the same age that Kimo is now)
x=4
..
How old are they now?
Jennifer's age=13
Kimo's age=17